OpenStudy (anonymous):

Which of the following are vertical asymptotes of the function y = 3cot(1/2x) - 4?

OpenStudy (anonymous):

A. x = 0 B. x = pi/2 C. x = +- 2pi D. x = 3pi

OpenStudy (e.mccormick):

What part os giving you trouble with this?

OpenStudy (anonymous):

i just cant solve it. hate trig.

OpenStudy (e.mccormick):

If you plan on taking calculus, learn to love trig. How far do you get before you get stuck?

OpenStudy (anonymous):

i dont know how to even start it.

OpenStudy (e.mccormick):

OK, well, do you know what a vertical asymtote is? Lets start there.

OpenStudy (anonymous):

no

OpenStudy (e.mccormick):

AH! OK, that is probably the real problem. A vertical asymptote is a line the range of the equation does not cross. So it goes from \(-\infty\) to \(\infty\) and never gets crossed. Where this happens with cotangent is when the bottom part is 0. Know how cotangent is cosign over sign? That is the bottom part I mean. When the sign part would be 0, the equation becomes undefined and there is a vertical asymptote.

OpenStudy (e.mccormick):

Well, sine... LOL. Not sign. Mistyped there, but you should get the point.

OpenStudy (anonymous):

so the answer is A?

OpenStudy (e.mccormick):

That is pne answer. This is probably multiple selectables. There is more than one answer.

OpenStudy (e.mccormick):

Test each of the points in say a calculator. Any place you get an error for an answer is it. That is one way. Knowing the unit circle is another.

OpenStudy (anonymous):

i dont have a calc left it at skool

OpenStudy (e.mccormick):

ah. Now, one question, is that \(\frac{1}{2}x\) or \(\frac{1}{2x}\) just to be sure I am looking at the question correctly.

OpenStudy (anonymous):

first one

OpenStudy (e.mccormick):

So we could call it x/2. So, any place \(\sin{\frac{x}{2}}=0\). If you look carefully, nothing else about that problem matters. The -4 shifts the graph down four, but does not change the period. The 3 stretches it, but does not change the period. So all that matters is when \(\sin{\frac{x}{2}}=0\).

OpenStudy (anonymous):

wait i m calculating

OpenStudy (e.mccormick):

np. If you have the unit circle, you can just try those four points and see what they become.

OpenStudy (e.mccormick):

Just remember it is \(\frac{x}{2}\). That DOES change the period!

OpenStudy (anonymous):

so how should i try solving this in the calulator

OpenStudy (e.mccormick):

Well, lets take a point. Say that \(\frac{\pi}{2}\). We put it into \(\frac{x}{2}\) That gives us:\[\frac{\frac{\pi}{2}}{2}\implies \frac{\pi}{4}\] So, is \(\sin(\frac{\pi}{4})=0\)?

OpenStudy (anonymous):

0.75....

OpenStudy (e.mccormick):

So that is NOT a vertical asymptote. Now do the same thing for the next point. So on and so forth. Any point that has a sine of 0 means the cotangent would be undefined and would be a vertical asymptote.

OpenStudy (anonymous):

so C and A?

OpenStudy (e.mccormick):

Yep. And here is a reference with your points and a few others graphed. The vertical asymptotes are where the cotangent gets close, but never crosses. https://www.desmos.com/calculator/swir7owcxw

OpenStudy (anonymous):

tysm

OpenStudy (e.mccormick):

I hope the graph makes it a little clearer what this is all about. Asymptotes, vertical and otherwise, become more important as you go on in math.

OpenStudy (e.mccormick):

On the graph, you can remove the extra lines I put in and just see the points you are concerned with. It was just an easy way to visualize it for you.

OpenStudy (anonymous):

ok thx

OpenStudy (e.mccormick):

Well, don't have too much fun!